If I have a normal distribution that has a mean of 100 and standard deviation of 20, and if I draw a sample of size n=40, how do I find the probability that all the 40 sampled values are less than 120? The 40 values being sampled are done so without replacement.
Answer by stanbon(75887)

If I have a normal distribution that has a mean of 100 and standard deviation of 20, and if I draw a sample of size n=40, how do I find the probability that all the 40 sampled values are less than 120? The 40 values being sampled are done so without replacement.
-----------
Find the z-value of 120:
z(120)=(120-100)/20 = 1
P(an individual in the population < 120) = P(-10< z< 1) = 0.8413..
--------------
P(40 selected at random < 120) = 0.8413...^40 = 0.00099761
=============
Cheers,
Stan H.